​Bondholders' expected rate of return​) The market price is ​$1000 for a 16​-year bond ​($1000 par​ value) that pays 8 percent interest ​(4 percent​ semiannually). What is the​ bond's expected rate of​ return? The​ bond's expected annual rate of return is
Answer #1

use financial calculator function as below

FV=1000 is the face value or par value of the bond

PV=-1000 is the present value or current price of the bond

N=16*2=32 semi annual period

PMT=1000*8%/2=40

Click CPT

Click I/Y=4.00% is semi annual rate

The​ bond's expected annual rate of return is=2*semi annual rate=2*4.00%=8.00% (answer)

the above is answer..